ganeshie8 (ganeshie8):

Here is a cool problem Find the voltage across each resistor as \(R\to\infty\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Applying KVL innocently gives \[10 - V_{r}-V_{r} = 0\] Rearranging gives \[V_R = 5\] Does this still hold as \(R\to \infty\) ?

ganeshie8 (ganeshie8):

I think two holes appear in the circuit as \(R\to\infty\)  What can you say about the voltage across each hole ?